OpenStudy (dtan5457):

If x+1/x=7, compute x^3+1/x^2

OpenStudy (dtan5457):

So far I have x^2+1=7x x^2=7x-1 x^3+1/7x-1 x^2x/(1)+1/(7x-1) when I cross out 7x-1 i get x+1. then what?

OpenStudy (dtan5457):

@Hero

hero (hero):

Hang on a min.

OpenStudy (dtan5457):

okay thanks

hero (hero):

( x+1)/x=7 multiply both sides by × x +1 =7x Subtract × from both sides 1=6x x= 1/6

hero (hero):

Plug × to other expression

OpenStudy (dtan5457):

wait so all my x^2+1=7x was incorrect?

hero (hero):

. Is the original problem you posted correct as written?

OpenStudy (dtan5457):

\[x+\frac{ 1 }{ x }=7, x^3+\frac{ 1 }{ x^2 }=?\]

hero (hero):

± see, Better you use later to post from now on.

hero (hero):

latex *

OpenStudy (dtan5457):

okay

OpenStudy (dtan5457):

so does your solution still stand?

OpenStudy (dtan5457):

@jdoe0001

hero (hero):

No, it doesn't

OpenStudy (dtan5457):

any ideas?

hero (hero):

Calculate ( x +1/ x ) ^ 2 and (x+1 / x ) ^ 3 )

OpenStudy (dtan5457):

i dont get it sorry

OpenStudy (dtan5457):

one of my friends even said its unsolvable without quadratic formula but im pretty sure my teacher doesnt want a quadratic use

hero (hero):

( x + 1 / x ) ^ 2 = 7^2

hero (hero):

If you don't get that I can't help you

hero (hero):

Expand the left side

OpenStudy (dtan5457):

\[x^2+\frac{ 1 }{ x^2 }+x+\frac{ 1 }{ x^2 }\]

hero (hero):

That's not what I got

OpenStudy (dtan5457):

lol i give up, i'll ask my teacher tomorrow. thanks

hero (hero):

You give up too easy.

OpenStudy (dtan5457):

Been working on 4 problems for about 2 hours, it's time too move on if I want decently sufficient sleep. I've been getting mixed answers and solutions so I'll just let my teacher explain.

OpenStudy (dtan5457):

it can be solved with the quadratic formula so I just used that. Not sure if that's what my teacher wants but i'll see.

hero (hero):

Nope, it probably isn't. There's another way to solve that your teacher may be looking for.